# Magnetic tunnel junction equivalent circuit modeling

### v2

@@ Abstract
This work presents an equivalent circuit model for Magnetic Tunnel Junctions (MTJs) that accurately captures their magnetization dynamics and electrical behavior. Implemented in LTspice, the model is validated against direct numerical solutions of the Landau-Lifshitz-Gilbert-Slonczewski (LLGS) equation. It effectively simulates essential spintronic phenomena, including ferromagnetic resonance, field- and spin-torque-induced switching, and spin-torque-induced oscillations. Simulation results demonstrate strong agreement between LTspice and LLGS solutions, confirming the model accuracy and utility for efficient circuit-level analysis of spintronic devices. The ability to incorporate time-dependent magnetic fields and voltage inputs makes the proposed model suitable for diverse applications such as neuromorphic computing, microwave signal processing, and spintronic memory technologies. By providing a computationally efficient yet physically accurate circuit representation, this work enables seamless integration of MTJs into larger electronic systems, potentially accelerating the development of advanced spintronic circuit architectures.
@@ 1 INTRODUCTION
In the past two decades, magnetic tunnel junctions (MTJs) revolutionized data storage by enabling high-density recording in disk-based drives [Chappert, 2007]. Beyond storage, MTJs integrated with CMOS technology have been commercially deployed in magnetoresistive random-access memory (MRAM) [Alzate 2015, Edelstein 2020], which utilizes non-volatility and fast switching properties of MTJs to provide energy-efficient and high-speed memory [Ikegawa 2020, Ikegawa 2021]. Beyond digital applications, MTJs can also be used as current-driven microwave sources [Fuchs 2004, Villard 2009], microwave detectors [Tulapurkar 2005], energy harvesters [Fang 2019] and more advanced microwave signal processing devices [Litvinenko 2022]. Recently, MTJs have also been explored neuromorphic computing applications [Romera 2018, Louis 2024, Rodrigues 2023].
@@
An MTJ consists of two ferromagnetic layers separated by a thin insulating barrier [Singh 2019, Yuasa 2007]. The operational principle of an MTJ is based on the tunneling magnetoresistance effect, which is sensitive to the relative orientation of magnetic moments in the ferromagnetic layers [Yuasa 2007], and on the spin-transfer-torque (STT) effect [Slonczewski 1989, Ralph 2008] that describes influence of electric current on magnetization dynamics. Advanced nanoscale fabrication methods have enabled integration of MTJs into complex electronic systems, expanding their use in memory and logic devices [Alzate 2015, Edelstein 2020]. These distinctive properties position MTJs as important components in next-generation memory, analog, and neuromorphic computing systems [Incorvia 2024].
@@
Figure 1: (a) Schematic of the MTJ, illustrating the fixed and free layers. (b) Equivalent circuit representation of the MTJ free layer.
Circuit simulation is essential in the design and analysis of electronic systems [Najm 2010]. An accurate circuit-equivalent model of an MTJ facilitates its integration into larger circuits and allows comprehensive performance evaluation. Several MTJ circuit models have been previously proposed [Xu 2012, Panagopoulos 2012, Kim 2015], but they are based on formal mathematical equivalence of magnetization and circuit equations and do not faithfully represent physical aspects of MTJ behavior. For example, they do not contain energy storing elements (capacitance or inductance) that represent magnetic energy of an MTJ, or resistive elements describing Gilbert damping mechanism.
@@
This letter introduces an equivalent circuit model for MTJs that accurately reflects their underlying physical characteristics. Within the proposed model, conservative magnetic interactions (e.g., effect of a bias magnetic field or magnetic anisotropy) are modeled by equivalent nonlinear capacitance and inductance. Dissipative magnetic torques are represented by equivalent resistance (Gilbert damping) and current and voltage sources (STT). The model is realized as a standalone element in LTspice electrical engineering software and its validity is demonstrated by comparing LTspice simulation results with direct numerical simulations of the Landau-Lifshits-Gilbert-Slonczewski (LLGS) equation.
@@ 2 Physics-Based Equivalent MTJ Circuit
An MTJ comprises a ferromagnetic fixed layer, a non-magnetic spacer, and a ferromagnetic free layer, as shown in Fig. 1(a). The magnetization of the fixed layer is represented by $\mathbf{p}$, while the free layer magnetization is denoted by $\mathbf{m}$. The term “free layer” indicates that the magnetization direction can be altered by external stimuli, such as an applied magnetic field $\mathbf{B}_{\mathrm{ext}}$ or electric current $I$. In contrast, the fixed layer has a stable magnetization direction that remains largely unaffected by external perturbations. The dynamic behavior of the free layer arises from the interplay between magnetic fields, damping effects, and spin-transfer torque, producing complex magnetization dynamics. These dynamics determine the electrical behavior and switching characteristics of MTJ-based devices [Slavin 2009].
When an electric current $I$ passes through the MTJ, it encounters a resistance denoted as $R_{\mathrm{mtj}}$. The resistance $R_{\mathrm{mtj}}$ reaches a minimum value, $R_{\mathrm{P}}$, when the magnetization vectors $\mathbf{p}$ and $\mathbf{m}$ are parallel. Conversely, the resistance is maximized, taking the value $R_{\mathrm{AP}}$, when $\mathbf{p}$ and $\mathbf{m}$ are antiparallel. If the vectors $\mathbf{p}$ and $\mathbf{m}$ are orthogonal, the resistance becomes $R_{\perp}$, defined by $R_{\perp}=2R_{\mathrm{AP}}R_{\mathrm{P}}/(R_{\mathrm{AP}}+R_{\mathrm{P}})$. A general expression for the resistance of an MTJ is provided by [Slonczewski 1989, Wang 2009],
$R_{\mathrm{mtj}}=\frac{R_{\perp}}{1+\eta^{2}(\mathbf{p}\cdot\mathbf{m})},$ (1)
where $\eta$ is the dimensionless spin polarization efficiency, given by $\eta=\sqrt{(R_{\mathrm{AP}}-R_{\mathrm{P}})/(R_{\mathrm{AP}}+R_{\mathrm{P}})}$.
The magnetization dynamics of the free layer can be effectively described using the macrospin approximation, which is governed by the Landau-Lifshitz-Gilbert-Slonczewski (LLGS) equation [Slavin 2009], given by
@@
where $\gamma$ is the gyromagnetic ratio, and $\alpha_{G}$ is the Gilbert damping parameter. The spin current $\mathbf{J}_{s}$ can be written as $\mathbf{J}_{s}=(\gamma\hbar\eta/2eM_{s}V_{s})(V_{\mathrm{mtj}}/R_{\perp})\mathbf{p}$, where $\hbar$ is the reduced Planck constant, $\mu_{0}$ is the vacuum magnetic permeability, $e$ is the elementary charge, $V_{s}$ represents the volume of the free layer, and $V_{\mathrm{mtj}}$ is the voltage across the MTJ. The effective magnetic field $\mathbf{B}_{\mathrm{eff}}$ is given by $\mathbf{B}_{\mathrm{eff}}=-(1/M_{s}V_{s})\partial E(\mathbf{m})/\partial\mathbf{m}$, where $E(\mathbf{m})$ is the macrospin energy which is given by
$E(\mathbf{m})=M_{s}V_{s}\Big{[}-B_{e}\cdot\mathbf{m}+\tfrac{1}{2}B_{d}(\mathbf{m}\cdot\hat{\mathbf{z}})^{2}-\tfrac{1}{2}B_{a}(\mathbf{m}\cdot\hat{\mathbf{x}})^{2}\Big{]},$ (3)
where $\mathbf{B}_{e}$ is the externally applied field, $B_{a}$ is the easy-axis anisotropy field in the MTJ plane, and $B_{d}$ is the effective demagnetizing field having contributions from dipolar field and perpendicular magnetic anisotropy (PMA).
As discussed above, Eq. (2) describes the free layer magnetization dynamics. Since MTJ resistance, $R_{\mathrm{mtj}}$, depends on the relative orientation of magnetization vectors, variations in $\mathbf{m}$ directly affect device electrical behavior. In Eq. (2), the left side represents conservative precessional motion. On the right side, the first term describes Gilbert damping torque, and the second represents Slonczewski spin-transfer torque.
The macrospin model Eq. (2) can be represented by an equivalent nonlinear LC circuit by defining effective capacitor charge $Q$ and inductor flux $\Phi$ as
$Q=\frac{e}{\hbar}\frac{M_{s}V_{ol}}{\gamma}\cos\theta,\hskip 11.38109pt\Phi=\frac{\hbar}{e}\phi,$ (4)
where $\theta$ and $\phi$ are, respectively, the polar and azimuthal angles of the magnetization vector $\mathbf{m}$ (see Fig. 1(a)). For detailed derivation of the model see [Louis 2025].
Figure 1(b) presents the equivalent circuit LC model with additional elements that capture all aspects of the dynamics of the free MTJ layer. In this representation, the capacitor $C$ and inductor $L$ correspond to the conservative precessional dynamics terms in the LLGS equation. The resistors $R_{C}$ and $R_{L}$ model the dissipation associated with the Gilbert damping mechanism. The current source $I_{s}$ and the voltage source $V_{s}$ arise from Slonczewski spin-transfer torque.
Following standard definitions for electrical circuits, the current through the capacitor $I_{C}$ and the voltage at the inductor $V_{L}$ are defined as
$I_{C}=\partialderivative{Q}{t},\hskip 11.38109ptV_{L}=\partialderivative{\Phi}{t}.$ (5)
Likewise, the voltage at the capacitor $V_{C}$ and the current through the inductor $I_{L}$ are related to the energy $E$ of the LC circuit by
@@
For an MTJ with energy function Eq. (3), these definitions give
@@
Here, $Z_{s}=\hbar^{2}\gamma/(eM_{s}V_{s})$ is a “characteristic impedance” of the macrospin – a quantity with dimension of electrical resistance that naturally appears in the equivalent circuit model.
These equations represent constitutive relations for equivalent macrospin capacitance and inductance and can be directly incorporated into modern electrical engineering software (e.g., LTspice) using various behavioral elements.
Other parameters of the equivalent circuit Fig. 1(b) are:
$R_{C}=\frac{\alpha_{G}Z_{s}}{\sin^{2}\theta},$ (9)
$R_{L}=\frac{Z_{s}}{\alpha_{G}\sin^{2}\theta},$ (10)
$I_{s}=\frac{\eta}{2}\Big{[}p_{z}\sin\theta-(p_{x}\cos\phi+p_{y}\sin\phi)\Big{]}\sin\theta\frac{V_{\textrm{mtj}}}{R_{\perp}},$ (11)
$V_{s}=\frac{\eta}{2}\Big{[}\frac{p_{x}\sin\phi-p_{y}\cos\phi}{\sin\theta}\Big{]}\frac{V_{\textrm{mtj}}}{R_{\perp}}Z_{s}.$ (12)
For a complete derivation of the equivalent circuit model see [Louis 2025].

### v3

@@ Abstract
This work presents an equivalent circuit model for magnetic tunnel junctions (MTJs) that accurately reproduces their magnetization dynamics and electrical behavior within the macrospin approximation. The model is validated through direct numerical simulations of the Landau–Lifshitz–Gilbert–Slonczewski (LLGS) equation, encompassing ferromagnetic resonance, field- and spin-torque-induced switching, and spin-torque-induced oscillations. Simulation results exhibit excellent agreement between the equivalent circuit model and the LLGS-based simulations, confirming the model accuracy and utility for efficient circuit-level analysis of MTJs. The capability of handling time-dependent magnetic fields and voltage-driven excitations renders the proposed model applicable to diverse areas, including neuromorphic computing, microwave signal processing, and spintronic memory technologies. By providing a computationally efficient yet physically rigorous circuit representation, this work facilitates seamless integration of MTJs into complex electronic systems, thereby accelerating the advancement of novel spintronic circuit architectures.
@@ 1 INTRODUCTION
In the past two decades, magnetic tunnel junctions (MTJs) have revolutionized data storage by enabling high-density recording in disk-based drives [Chappert, 2007]. MTJs have also been integrated with CMOS technology and commercialized in magnetoresistive random-access memory (MRAM) [Alzate 2015, Edelstein 2020] which leverages the non-volatility and fast switching properties of MTJs for energy-efficient, high-speed memory [Ikegawa 2020]. Beyond digital applications, MTJs may also be used as microwave sources [Fuchs 2004, Villard 2009], microwave detectors [Tulapurkar 2005], energy harvesters [Fang 2019], advanced microwave signal processing devices [Litvinenko 2022], and even in neuromorphic computing [Romera 2018, Louis 2024, Rodrigues 2023].
@@
An MTJ consists of two ferromagnetic layers separated by a thin insulating barrier [Singh 2019, Yuasa 2007]. The operational principle of an MTJ is based on the tunneling magnetoresistance effect, which is sensitive to the relative orientation of magnetic moments in the ferromagnetic layers [Yuasa 2007], and on the spin-transfer-torque (STT) effect [Slonczewski 1989, Ralph 2008] that describes influence of electric current on magnetization dynamics. Advanced nanoscale fabrication enabled integration of MTJs into complex electronic systems, expanding their use in memory and logic devices [Alzate 2015, Edelstein 2020]. These distinctive properties position MTJs as important components in next-generation analog and neuromorphic computing systems [Incorvia 2024].
@@
Figure 1: (a) Diagram of the MTJ, illustrating the free and fixed layers. (b) Equivalent circuit of the MTJ, as an intrinsic variable resistor. (c) Equivalent circuit representation of the MTJ free layer magnetization. (d) The complete MTJ circuit implemented in LTspice.
Circuit simulation is essential in the design and analysis of electronic systems [Najm 2010]. An accurate circuit-equivalent model of an MTJ facilitates its integration into larger circuits and allows comprehensive performance evaluation, while preserving insight into magnetization dynamics. Prior works [Xu 2012, Panagopoulos 2012, Kim 2015] have used LTspice and other circuit simulators to numerically solve the LLGS equation, for example by mapping angular variables onto circuit elements. However, these approaches do not establish a physical correspondence between the circuit elements and the underlying energy storage, dissipation, or torque mechanisms present in the MTJ. There are models [Bendjeddou 2023, Abbasi 2017] that can emulate the MTJ impedance at high frequency, but do not take into account MTJ behavior under the influence of static inputs.
@@
This letter introduces an equivalent circuit model for MTJs that accurately reflects their underlying physical characteristics. Within the proposed model, conservative magnetic interactions (e.g., effect of a bias magnetic field or magnetic anisotropy) are modeled by equivalent nonlinear capacitance and inductance. Dissipative magnetic torques are represented by equivalent resistance (Gilbert damping) and current and voltage sources (STT). The model is realized as a standalone element in LTspice circuit simulation software and its validity is demonstrated by comparing LTspice simulation results with direct numerical simulations of the Landau-Lifshits-Gilbert-Slonczewski (LLGS) equation.
@@ 2 Physics-Based Equivalent MTJ Circuit
An MTJ comprises a ferromagnetic fixed layer, a non-magnetic spacer, and a ferromagnetic free layer, as shown in Fig. 1(a). The magnetization of the fixed layer is represented by $\mathbf{p}$, while the free layer magnetization is denoted by $\mathbf{m}$. The term “free layer” indicates that the magnetization direction can be altered by external stimuli, such as an applied magnetic field $\mathbf{B}_{\mathrm{ext}}$ or electric current $I$. In contrast, the fixed layer has a stable magnetization direction that is largely unaffected by external perturbations. The dynamic behavior of $\mathbf{m}$ arises from the interplay between magnetic fields, damping effects, and spin-transfer torque, producing complex magnetization dynamics. These dynamics determine the electrical behavior and switching characteristics of MTJ-based devices [Slavin 2009].
When an electric current $I$ passes through the MTJ, it encounters a resistance denoted $R_{\mathrm{mtj}}$. The resistance $R_{\mathrm{mtj}}$ reaches a minimum value, $R_{\mathrm{P}}$, when the magnetization vectors $\mathbf{p}$ and $\mathbf{m}$ are parallel. Conversely, the resistance is maximized, taking the value $R_{\mathrm{AP}}$, when $\mathbf{p}$ and $\mathbf{m}$ are antiparallel. If the vectors $\mathbf{p}$ and $\mathbf{m}$ are orthogonal, the resistance becomes $R_{\perp}$, defined by $R_{\perp}=2R_{\mathrm{AP}}R_{\mathrm{P}}/(R_{\mathrm{AP}}+R_{\mathrm{P}})$. A general expression for the resistance of an MTJ is [Slonczewski 1989, Wang 2009],
$R_{\mathrm{mtj}}(\mathbf{m})=\frac{R_{\perp}}{1+\eta^{2}(\mathbf{p}\cdot\mathbf{m})},$ (1)
where $\eta$ is the dimensionless spin polarization efficiency, given by $\eta=\sqrt{(R_{\mathrm{AP}}-R_{\mathrm{P}})/(R_{\mathrm{AP}}+R_{\mathrm{P}})}$. The MTJ shown in Fig. 1(a) can be represented by the circuit in Fig. 1(b), which incorporates an intrinsic magnetoresistor symbol as defined in [IEEE Standard] 2.1.7. The resistance $R_{\mathrm{mtj}}(\mathbf{m})$ is intrinsically nonlinear, depending on the free layer magnetization $\mathbf{m}$.
The magnetization dynamics of the free layer can be effectively described using the macrospin approximation, which can be modeled by the Landau-Lifshitz-Gilbert-Slonczewski (LLGS) equation [Slavin 2009], given by
@@
where $\gamma$ is the gyromagnetic ratio, and $\alpha_{G}$ is the Gilbert damping parameter. The spin current $\mathbf{J}_{s}$ can be written as $\mathbf{J}_{s}=(\gamma\hbar\eta/2eM_{s}V_{o})(V_{\mathrm{mtj}}/R_{\perp})\mathbf{p}$, where $\hbar$ is the reduced Planck constant, $\mu_{0}$ is the vacuum magnetic permeability, $e$ is the elementary charge, $V_{o}$ represents the volume of the free layer, and $V_{\mathrm{mtj}}$ is the voltage across the MTJ. The effective magnetic field $\mathbf{B}_{\mathrm{eff}}$ is given by $\mathbf{B}_{\mathrm{eff}}=-(1/M_{s}V_{o})\partial E(\mathbf{m})/\partial\mathbf{m}$, where $E(\mathbf{m})$ is the macrospin energy which is given by
$E(\mathbf{m})=M_{s}V_{o}\Big{[}-\mathbf{B}_{e}\cdot\mathbf{m}+\tfrac{1}{2}B_{d}(\mathbf{m}\cdot\hat{\mathbf{z}})^{2}-\tfrac{1}{2}B_{a}(\mathbf{m}\cdot\hat{\mathbf{x}})^{2}\Big{]},$ (3)
where $\mathbf{B}_{e}$ is the externally applied field, $B_{a}$ is the easy-axis anisotropy field in the MTJ plane, and $B_{d}$ is the effective demagnetizing field having contributions from dipolar field and perpendicular magnetic anisotropy (PMA). Please note that left side of (2) represents conservative precession, while the right side includes Gilbert damping torque and Slonczewski spin-transfer torque (STT).
Equation (2) models the magnetization dynamics of the free layer and can be used to determine the macrospin vector $\mathbf{m}(\theta,\phi)$ as a function of time, where $\theta$ and $\phi$ are the polar and azimuthal angles of $\mathbf{m}$ (see Fig. 1(a)). This dynamic behavior can be equivalently modeled by a nonlinear LC circuit by defining an effective capacitor charge $Q$ and inductor flux $\Phi$
$Q=\frac{e}{\hbar}\frac{M_{s}V_{o}}{\gamma}\cos\theta,\hskip 11.38109pt\Phi=\frac{\hbar}{e}\phi.$ (4)
The voltage $V_{C}$ across a capacitor and the current $I_{L}$ through an inductor can be related to the energy $E$ of an LC circuit by
@@
For an MTJ with an energy function (3), these definitions give
@@
Here, $Z_{s}=\hbar^{2}\gamma/(eM_{s}V_{o})$ is a “characteristic impedance” of the macrospin. Equations (6) and (7) represent constitutive relations for equivalent macrospin capacitance and inductance; that is, they define how $V_{C}$ and $I_{L}$ relate directly to the magnetic state variables ($\theta$ and $\phi$), thereby providing the bridge between magnetization dynamics and circuit-level electrical behavior.
From the constitutive relations (6) and (7), a capacitance $C$ can determined as $C=\left(\partial^{2}E/\partial Q^{2}\right)^{-1}$, or
$C=\frac{1}{\gamma Z_{s}}\Big{[}\frac{1}{B_{d}+B_{a}\cos^{2}\phi+(B_{e,x}\cos\phi+B_{e,y}\sin\phi)/\sin^{3}\theta}\Big{]}.$ (8)
Likewise, an inductance $L=\left(\partial^{2}E/\partial\Phi^{2}\right)^{-1}$ can be determined,
$L=\frac{Z_{s}}{\gamma}\Big{[}\frac{1}{(B_{e,x}\cos\phi+B_{e,y}\sin\phi)\sin\theta+B_{a}\cos 2\phi\sin^{2}\theta}\Big{]},$ (9)
Please note that both the capacitance and inductance depend explicitly on the external magnetic field and the free layer magnetization angles $\theta$ and $\phi$. For notational simplicity, we will express them as functions of the free layer magnetization: $C\rightarrow C(\mathbf{m})$ and $L\rightarrow L(\mathbf{m})$.
Fig. 1(c) presents the equivalent circuit model that incorporates $C(\mathbf{m})$ and $L(\mathbf{m})$ along with additional elements:
$R_{C}=\frac{\alpha_{G}Z_{s}}{\sin^{2}\theta},\hskip 11.38109ptR_{L}=\frac{Z_{s}}{\alpha_{G}\sin^{2}\theta},$ (10)
$I_{s}(V_{\textrm{mtj}})=\frac{\eta}{2}\Big{[}p_{z}\sin\theta-(p_{x}\cos\phi+p_{y}\sin\phi)\Big{]}\sin\theta\frac{V_{\textrm{mtj}}}{R_{\perp}},$ (11)
$V_{s}(V_{\textrm{mtj}})=\frac{\eta}{2}\Big{[}\frac{p_{x}\sin\phi-p_{y}\cos\phi}{\sin\theta}\Big{]}\frac{V_{\textrm{mtj}}}{R_{\perp}}Z_{s}.$ (12)
A complete derivation of these expressions and full development of the equivalent circuit (Fig. 1(c)) is provided in [Louis 2025].
This equivalent circuit accurately captures the complete dynamics of $\mathbf{m}$. In direct correspondence with (2), the capacitor $C(\mathbf{m})$ and inductor $L(\mathbf{m})$ model conservative precession, resistors $R_{C}$ and $R_{L}$ represent Gilbert damping, and the sources $I_{s}(V_{\textrm{mtj}})$ and $V_{s}(V_{\textrm{mtj}})$ include MTJ voltage dependence, capturing STT effects.
This paragraph outlines the procedure for simulating the MTJ dynamics using the equivalent circuit model. The simulation begins by applying the inputs $V_{\textrm{mtj}}$ and $I$ to the charge circuit in Fig. 1(b), while specifying the external magnetic field $\mathbf{B}_{e}$. These inputs drive the spin circuit shown in Fig. 1(c), which outputs the capacitor current $I_{C}$ and the inductor voltage $V_{L}$. According to standard circuit definitions, these are related to the state variables by
[…]
From $I_{C}$ and $V_{L}$, the circuit determines $Q$ and $\Phi$, which are then used in Eq. (4) to compute the free layer magnetization angles $\theta$ and $\phi$. This yields the magnetization vector $\mathbf{m}(\theta,\phi)$, allowing the MTJ resistance $R_{\mathrm{mtj}}(\mathbf{m})$ to be evaluated through Eq. (1). Finally, this resistance is fed back into the charge circuit in Fig. 1(b), completing one self-consistent iteration of the simulation.
The circuits shown in Fig. 1(b) and Fig. 1(c) were implemented in LTspice, a free circuit simulation software provided by [Analog Devices]. To simplify simulation workflows, these circuits were encapsulated into a single LTspice component, illustrated in Fig. 1(d). This standalone component features input and output terminals labeled “$+$” and “--” for electrical signals, along with additional input ports for specifying the external magnetic field $\mathbf{B}_{e}$. The output ports $\theta$ and $\phi$ can be used to monitor internal state of the MTJ. The complete LTspice model, including simulation examples and documentation, is available at [MTJ Spice Models].
The equivalent circuit framework is designed to be modular and easily extendable. For instance, incorporating additional physical effects such as field-like torque involves modifying the macrospin energy expression in (3), which in turn alters the constitutive relations in (6) and (7). Similarly, configurations with more than one fixed magnetic layer, as explored in [Louis2024], can be modeled by extending Eq. (1) to include multiple spin polarization terms and updating the spin circuit in Fig. 1(c) with additional dependent sources. These examples highlight the flexibility of the model, making it well-suited for simulating a wide range of spintronic device architectures. These extensions follow naturally from the underlying energy-based formulation, and a comprehensive derivation of the equivalent circuit model is provided in [Louis 2025].
Although this work focuses on a single macrospin, spatially non-uniform phenomena, such as spin-wave excitations and domain formation, could be approximated by networks of coupled macrospin circuit elements. This generalization offers a promising route to incorporate micromagnetic effects into circuit-level simulations.